OpenStudy (anonymous):

25 is what percent of 23? write the proportion too please!

OpenStudy (beccaboo333):

Do you need a full number or percentage?

OpenStudy (razzanide):

23/25 = 92/100 = 92% And I believe the proportion is just 23/25? I might be wrong about that bit though.